OpenStudy (anonymous):

x + xy + xyz = If x =1, y = 10, and z = 1/5

OpenStudy (4everaddicted2anime):

This comes out to \[1+1(10)+1(10)(\frac{ 1 }{ 5 })=?\]

OpenStudy (anonymous):

yeah but really its 1 + 1 x 10 + 1 x 10 x 1/5=

OpenStudy (4everaddicted2anime):

That is what I just said

OpenStudy (anonymous):

i know I was just putting multiplication sembles instead of perinthesis

OpenStudy (4everaddicted2anime):

Okay. If you simplify the parts where you are multiplying 1 and a number you should get \[1+10+10(\frac{ 1 }{ 5 })=?\]

OpenStudy (4everaddicted2anime):

When you simplify 10(1/5) you hsould get 2. Now you just need to add 1+10+2 to get your answer

OpenStudy (anonymous):

ok thanks

OpenStudy (4everaddicted2anime):

yw